A wooden board 41 feet long is a cut into three pieces. The 3rd piece is 8ft longer than the 1st piece and the 2nd piece is 5 less than the longer piece. Find the length of the three pieces. so that the length of each piece add up to 4
Answer by stanbon(75887)

A wooden board 41 feet long is a cut into three pieces.
---------
The 3rd piece is 8ft longer than the 1st piece and the 2nd piece is 5 less than the longer piece.
Let length of the 1st piece be "x".
Then the 3rd piece has length "x+8".
And the 2nd piece has length "x+8-5" = x+3
-----------------------------------------------------
Find the length of the three pieces. so that the length of each piece
add up to 41.
----------------------------------
Equation:
x + x+8 + x+3 = 41
3x + 11 = 41
3x = 30
x = 10 ft. (length of the 1st ft.)
x+3 = 13 ft. (length of the 2nd ft.)
x+8 = 18 ft. (length of the 3rd ft.)
=======================================
Cheers,
Stan H.